OpenStudy (anonymous):

Thevolumeincubicfeetofaboxcanbeexpressedas (푥)=푥ﰆ −6푥ﰇ +8푥,or as the product of three linear factors with integer coefficients. The width of the box is 2 – x. Factor the polynomial to find linear expressions for the height and the width.

OpenStudy (anonymous):

\[(x)=x^3-6x^2+8x\]

OpenStudy (campbell_st):

well start by looking at \[f(x) = x(x^2 -6x + 8)\] so i linear factor is x factor the quadratic in brackets to find the other linear factors.

OpenStudy (anonymous):

let me be honest , ihave no clue ive been learing off the internet but im completly lost w/ this

OpenStudy (campbell_st):

ok... so look for 2 numbers that multiply to 8 and add to -6, they are both negative.

OpenStudy (campbell_st):

well you know the width 2 - x and there is a factor of x so the 3rd factor is (4 - x) so your volume is f(x) = x(2-x)(4 -x) so if width is (2 - x) height is x length is (4 - x)

OpenStudy (anonymous):

ook im getting it thank you!

OpenStudy (anonymous):

how do i put in linear expression @campbell_st ?

OpenStudy (anonymous):

tahts weird i used to live on campbell st in chicago

OpenStudy (campbell_st):

the linear expressions are just x, 2 -x and 4 -x linear just means the highest power is 1 lol... well thats weird... I lived on a campbell_st in Australia.

OpenStudy (campbell_st):

and an expression is just an arrangement of letters and numbers. if you include an equals sign, you get an equation.

OpenStudy (anonymous):

ugh for some reason im really slow on this and not putting in right equation

OpenStudy (campbell_st):

well the 3 things multiplied given the volume V = length x width x height the 3 linear expressions you have are length = x width = 2 -x so height = 4 - x length and height are interchangable.

OpenStudy (anonymous):

oh wow gotcha ... #embarrasing